# NuHepMC: A standardized event record format for neutrino event generators Steven Gardiner1, Joshua Isaacson1 and Luke Pickering2

### v2

@@ A NuHepMC HepMC3::GenRunInfo instance must contain a
It is suggested that a NuHepMC HepMC3::GenRunInfo instance should contain all information required to reproduce the events in the vector. This may be stored in attributes with names beginning with "NuHepMC.Provenance". The information required will necessarily be generator-specific, but we suggest two attributes that would be helpful to downstream users:-
type: HepMC3::LongAttribute,
name: "NuHepMC.Provenance.NEvents -
type: Implementation defined,
name: "NuHepMC.Provenance.RNGState"-
This might be a single number used as the seed to initialize the random number generator (RNG). It could also be a more complicated description of the RNG state.
@@
While G.R.5 and G.R.6 explicitly do not require implementations to emit metadata for standard status codes defined in the HepMC3 standard, it is suggested that the complete list of status codes used by an implementation are included in the "NuHepMC.VertexStatusInfo" and "NuHepMC.ParticleStatusInfo" attributes.
@@ 2.4 Event Data
The event is used to store information about the event as a whole. An event is described by arbitrary metadata and a graph of particles (edges) and vertices (nodes), each with their own arbitrary metadata. The NuHepMC specifications for events are as follows:
@@
The HepMC3 standard places some constraints on valid event graphs, these constraints must be respected by valid NuHepMC events as we require full compatibility with HepMC3. More details of these constraints can be found in Ref. [22].Existing neutrino event generators often rely on effective descriptions of the nuclear environment in a neutrino–nucleus hard scattering process. This means that four-momentum is often not explicitly conserved for the neutrino–nucleus system. Energy and momenta can be exchanged with a nuclear remnant, which is not directly involved in a neutrino–nucleon hard scatter, through initial and final state interactions. Implementations are free to conserve four momentum and emit all physical initial and final state particles, including the fully-simulated nuclear remnant, but for those implementations where such a requirement is not feasible or would delay the adoption of this standard, P.C.2 reserves a non-standard particle number that can be used to represent a nuclear remnant that is not precisely simulated.
@@
The process ID for the primary physics process that is represented in the HepMC3::GenEvent must be recorded in a HepMC3::IntAttribute named "signal_process_id". The metadata defining this process ID must be stored according to G.R.4. -
@@
HepMC3::VectorDoubleAttribute, named "LabPos", with the same units as used when implementing E.R.4. See E.C.5 for how to optionally store time in this attribute. If the simulation did not involve a macroscopic geometry, then this variable may be set to [0, 0, 0]. -
@@
An event must contain exactly one particle with the incoming beam particle status code and one particle with the target particle status code (see P.R.1). We recommend that, in cases where the colliding initial-state particles are distinct, the more massive of the two should be considered the target. For neutrino scattering, the target will thus often be a complex nucleus or a free nucleon. In the case of equally massive particles, the choice to label one of them as the target is arbitrary.P.C.1 provides a convention for marking a constituent bound nucleon struck by the incoming beam particle in the event graph.
@@
All simulated incoming and outgoing physical particles must be written to the event. The storage of intermediate particles is considered an implementation detail.
@@
It is not appropriate to mandate a specific set of interaction processes and assign them IDs in this standard. Different models make different choices, and it is impossible to foresee modeling developments that would require new process IDs to be defined in the future. Instead, the ranges of IDs given below are recommended for high-level categorization of processes. Even if an implementation uses the convention in Table 1, it must still adhere to G.R.4.Identifier Process 100-199 Low-Energy Nuclear Scattering 200-299 Quasielastic 300-399 Meson Exchange Current 400-499 Resonance production 500-599 Shallow inelastic scattering 600-699 Deep inelastic scattering 700- Other process types Table 1: Process ID ranges for various process categories. Charged-current (CC) processes should have identifiers in the X00-X49 block, and neutral-current (NC) processes should have them in the X50-X99 block. Negative process IDs may be reserved for electromagnetic interactions in neutrino event generators that include them.
@@
The total cross-section for the incoming beam particle, with its specific energy, to interact with the target particle should be stored in a HepMC3::DoubleAttribute on the HepMC3::GenEvent, named "TotXS". See G.C.4 for conventions on signalling cross section units. -
@@
The total cross-section for the selected process ID for the incoming beam particle, with its specific energy, to interact with the target particle should be stored in a HepMC3::DoubleAttribute on the HepMC3::GenEvent, named "ProcXS". See G.C.4 for conventions on signalling cross section units. -
@@
Some simulations build up an estimate of the flux-averaged total cross section as they run. This makes implementing G.C.5 impractical in many cases. As an alternative, the built-in attribute HepMC3::GenCrossSection, accessed via GenEvent::cross_section should be used to store the current estimate of . A user can then use the best estimate provided with the last generated event to correctly scale an event rate to a cross-section prediction.For event generators that do not currently provide the value of in the output, B provides suggestions for algorithms for computing a running estimate and associated Monte Carlo statistical uncertainty as events are produced.
@@
If the "LabPos" attribute vector contains three entries then it is assumed to only contain the spatial position of the event. If it contains four entries, then the fourth entry is interpreted as the time of the event in seconds.
@@
For more complex beam simulations that can not adequately be described by a single energy or energy histogram (see G.C.7), it is suggested that the full parent decay history is included in the HepMC3::GenEvent. A full set of conventions for the description of beam particle production and parent particle decay chains (for the case of neutrino beams) is currently outside the scope of this specification, but generator implementations can signal that they adhere to this suggestion to notify users that some or all of the beam particle production information is included in the event.
@@ 2.5 Vertex Information
The vertices in a HepMC3 event are used to connect groups of incoming and outgoing particles. For the vertex information, there is only one requirement in the present version of the NuHepMC standard.
@@
We extend the HepMC3 definition of HepMC3::GenVertex::status to include the concept of a primary vertex, corresponding to the primary process (i.e., the one labelled according to E.C.1), and a final state interaction (FSI) summary vertex. The full set of defined status codes can be found in Table 2. Implementations are free to define specific vertex status codes to refer to individual FSI (or ISI) processes and output as much information as they require. However, a single summary vertex may be useful for some purposes if the full FSI history is very detailed or not often needed by users.Status Code Meaning Usage 0 Not defined Do not use 1 Primary interaction vertex Recommended for all cases 2 FSI Summary vertex Recommended for all cases 3-20 Reserved for future NuHepMC standards Do not use 21-999 Generator-dependent For generator usage Table 2: Set of vertex status codes Any secondary vertex included within a NuHepMC event may have a status between 21 and 999. Note that G.R.5 requires that all generator-specific status codes must be fully described by attributes stored in the HepMC3::GenRunInfo.
@@
When an interaction with a nucleon bound within a nucleus with definite kinematics is simulated, a HepMC3::GenVertex corresponding to the separation of the struck nucleon and the nuclear remnant may be included and assigned status code 21. If this convention is signalled via the mechanism described in G.C.1, then status code 21 need not be included in the implementation of G.R.5.
@@ 2.6 Particle Information
We extend the HepMC3 definition of HepMC3::GenParticle::status slightly to include the concept of a target particle. For neutrino scattering, this will usually be a target nucleus. The status codes are defined in Table 3.Status Code Description Usage 0 Not defined Do not use 1 Undecayed physical particle Recommended for all cases 2 Decayed physical particle Recommended for all cases 3 Documentation line Used for in/out particles in the primary process 4 Incoming beam particle Recommended for all cases 5-10 Reserved for future HepMC3 standards Do not use 11-19 Reserved for future NuHepMC standards Do not use 20 Target particle Recommended for all cases 21-200 Generator-dependent For generator usage 201- Simulation-dependent For simulation software usage Table 3: Particle status codes Note especially that any incoming real particle must have a status code of 4 or 20, and any outgoing real particle must have a status code of 1. This allows users to know at a glance which simulated particles must be considered “observable” and which are “internal” details of the calculation. Special care must be taken when including the effects of initial-state and final-state interactions.
Any internal particle included within a NuHepMC event may have a status in the range than 21-200. Note that G.R.6 requires that all generator-specific status codes must be fully described by attributes stored in the HepMC3::GenRunInfo.
@@
When an interaction with a bound nucleon with definite kinematics is simulated, the internal HepMC3::GenParticle corresponding to the bound nucleon should have status code 21. If this convention is signalled via the mechanism described in G.C.1, then status code 21 need not be included in the implementation of G.R.6. -
@@
HepMC3 places restrictions on all external particles in the event graph to facilitate automatic checking of four momentum conservation at the event graph level. As a result, we define the new particle number 2009900000 to correspond to a nuclear remnant pseudo-particle. This particle corresponds to an implementation detail that should be used to abide by HepMC3 constraints to not have external vertices, but should not be considered for physics analyses or onward simulation. The number is chosen, according to the PDG scheme [25, Sec. 45, p. 733], to be outside the range reserved for nuclear and quark-content particles and signals that it is a non-standard code by having the 6th and 7th least significant digits set to 9.If this convention is signalled via the mechanism described in G.C.1, then particle number 2009900000 need not be included in the implementation of G.R.8.
If the nuclear particle number of the remnant is known, it can be added as a HepMC3::IntAttribute on the HepMC3::GenParticle named
"remnant_particle_number".
@@ 3 Flux-averaged total cross section
Section 3.1 derives an expression for that is generally applicable to most simulations of interest for neutrino experiments, including those with a time-dependent neutrino source and a full treatment of the detector geometry. Example methods for obtaining Monte Carlo estimators of suitable for storage via E.C.4 are described in B. Section 3.2 describes some simple cases in which may be calculated directly, making them good candidates for situations in which event generators may implement G.C.5.
@@ 3.1 Derivation
is the flux-averaged total cross section. For a discussion on how these might be calculated using Monte-Carlo methods see B.
@@ 3.2 Analytic calculation under simplifying assumptions
In certain simple cases, the flux-averaged total cross section is known at the start of the run and can be computed analytically to implement G.C.5. A common example is the case of a point target of type that is exposed to uni-directional beam of (anti)neutrinos of species with a fixed emission time. In this case, we may write the differential flux and target density as
@@ 4 NuHepMC Generator Predictions
Using preliminary tools for converting proprietary neutrino event formats to the NuHepMC standard defined above, we demonstrate the utility of the common format with some cross-section predictions.
A comparison to a recent neutrino-argon cross-section measurement from the MicroBooNE collaboration [28] is shown on the left-hand side of Fig. 1. This comparison was made with the NUISANCE framework, which before this implementation of NuHepMC would have to have been built against GENIE, NEUT, and NuWro binaries of compatible versions to be able to generate the predictions shown in the figure. In the present workflow, results can be obtained using only generator-agnostic tools for interpreting events stored in the NuHepMC format. This relatively small reduction of technical requirements will dramatically lower the barrier to making high-quality prediction-data comparisons for many experts and non-experts in the neutrino community.
It is particularly worth noting that the different generator predictions in this comparison provide critical cross-section scaling information in different ways. Thanks to the provisions in this standard, NUISANCE is able to use G.C.1 to determine what scaling methods are available and then automatically scale the selected event rate to a cross-section measurement prediction without any user input or generator-specific code or knowledge. For reference, the NEUT and NuWro selected event projections are scaled via G.C.5, while GiBUU, GENIE, and Achilles are scaled via E.C.4.
We want to emphasise that the comparison to data is solely a demonstration of the prediction pipeline and not meant as a physics statement or the best / most recent predictions from each of the generators for the given data set.
@@
where is the flux-averaged total cross section (see Sec. 3), is the statistical weight of the -th event, is the number of generated events, and is the width of the -th bin. The symbol evaluates to unity when the value of from the -th event falls within the -th bin and zero otherwise. In the case of unit weight events, the value of for all events, and . One thus recovers the traditional definition of these estimators for unweighted events.
@@
where we have assumed that is exactly known. For cases where it is estimated using Monte Carlo techniques rather than directly calculated, the statistical uncertainty discussed in B also applies.

### v3

@@ 2.3.3 Suggestions
It is suggested that a NuHepMC HepMC3::GenRunInfo instance should contain all information required to reproduce the events in the vector. This may be stored in attributes with names beginning with "NuHepMC.Provenance". The information required will necessarily be generator-specific, but it should be noted that storing the full initial state of any random number generators used during the generator run is crucial for exact reproduction. -
@@
While G.R.9 and G.R.10 explicitly do not require implementations to emit metadata for standard status codes defined in the HepMC3 standard, it is suggested that the complete list of status codes used by an implementation are included in the "NuHepMC.VertexStatusInfo" and "NuHepMC.ParticleStatusInfo" attributes. -
G.S.3
Target Atomic Abundances:
Using the flux-averaged total cross section , it is straightforward to compute partial cross sections for specific final states, interaction modes, and components of the overall target material used in a run. An example of the last of these is a flux-averaged cross section for interactions with C when a composite CH target was simulated.In some cases, it can be desirable to use an existing vector to create cross-section predictions for a target material with a different composition than the one initially used, e.g., distributions for from a vector generated for a CH target. Making such predictions requires a knowledge of the flux-averaged relative abundances of the nuclides present in the original target material. For situations in which these are known at the start of a run, a HepMC3::VectorStringAttribute named "NuHepMC.TargetMaterialRelativeAtomicAbundance" may be stored in the HepMC3::GenRunInfo. Each element of the vector should be a string of the form "100ZZZAAA0[x]" containing a nuclear PDG code and the corresponding relative abundance enclosed in square brackets []. For nuclides with a relative abundance of , the nuclear PDG code may optionally be used alone. For example, a pure target might be represented in this format as ["1000060120[0.5]", "1000010010"].
For realistic experimental simulations that involve a detector and a non-uniform neutrino beam, the flux-averaged relative abundances are typically difficult to calculate analytically. While running estimates of the abundances can in principle be obtained as the generator runs (using techniques akin to those in E.C.4), alternative approaches will generally be simpler to implement. Cases in which a particular detector material is of interest can be addressed by selecting events based on their position within the lab frame (c.f. E.R.5). Further exploration of this topic is left to future standardization efforts focused on neutrino fluxes and detector geometries.
@@ 2.4 Event Data
The event is used to store information about one primary interaction process and any relevant secondary processes. An event is described by arbitrary metadata and a graph of particles (edges) and vertices (nodes), each with their own arbitrary metadata. The NuHepMC specifications for events are as follows:
2.4.1 Requirements
@@
The HepMC3 standard places some constraints on valid event graphs. These constraints must be respected by valid NuHepMC events as full compatibility with HepMC3 is required. More details of these constraints can be found in Ref. [23].Existing neutrino event generators often rely on effective descriptions of the nuclear environment in a neutrino–nucleus hard scattering process. This means that four-momentum may not be explicitly conserved for the neutrino–nucleus system. Energy and momentum can be exchanged with a nuclear remnant, which is often not treated carefully in the simulation of a primary neutrino–nucleon collision, via initial- and final-state interactions. Implementations are of course free to thoroughly model all involved particles, conserving four-momentum over the whole system, including a fully-simulated nuclear remnant. For those implementations where such a requirement is not feasible or would delay the adoption of this standard, P.C.2 reserves a non-standard particle number that can be used to represent a nuclear remnant that is not precisely simulated.
@@
The process ID for the primary physics process that is represented in the
HepMC3::GenEvent must be recorded in a HepMC3::IntAttribute named "signal_process_id". The metadata describing this process ID must be stored according to G.R.8. -
@@
HepMC3::VectorDoubleAttribute, named "lab_pos", with the same units as used when implementing E.R.4. See E.C.5 for how to optionally store time in this attribute. If the simulation did not involve a macroscopic geometry, then this variable may be set to [0, 0, 0]. -
@@
An event must contain exactly one particle with the incoming beam particle status code and one particle with the target particle status code (see P.R.1). It is recommended that, in cases where the colliding initial-state particles are distinct, the more massive of the two should be considered the target. For neutrino scattering, the target will thus often be a complex nucleus or a free nucleon. In the case of equally massive particles, the choice to label one of them as the target is arbitrary.n.b. A nucleon (or quark) bound within a nucleus should never be marked as the target particle; the nucleus itself should be considered the target. P.C.1 provides a convention for marking a constituent bound nucleon struck by the incoming beam particle in the event graph.
@@
All simulated incoming and outgoing physical particles must be written to the event. The storage of intermediate particles is considered an implementation detail. -
E.R.9
Event Frame of Reference:
As the primary use of HepMC3 is for experiment simulation, all particle momenta specified via standard HepMC3 mechanisms must be defined in the lab frame. From the information required by a standard HepMC3 event, arbitrary frame transformations should be possible and can be performed by the
HepMC3::GenEvent::boost method. If storing kinematic information in other frames of reference facilitates later calculations, this should be done via simulation-specific HepMC3::Attributes on the relevant HepMC3::GenParticles.
2.4.2 Conventions
@@
It is not appropriate to mandate a specific set of interaction processes and assign them IDs in this standard. Different models make different choices, and it is impossible to foresee modeling developments that would require new process IDs to be defined in the future. Instead, the ranges of IDs given below are provided as a categorization of processes to facilitate high-level analysis. If a given process naturally fits in one of these categories it can be helpful to choose IDs that follow this convention, but an explicit aim of this standard is to avoid constraining or defining what processes can be stored and described. n.b. Even if an implementation uses the convention in Table 1, it must still adhere to G.R.8.Identifier Process 100-199 Low-energy nuclear scattering 200-299 Quasi-elastic 300-399 Multi-body scattering 400-499 Baryon resonance production 500-599 Shallow inelastic scattering 600-699 Deep inelastic scattering 700- Other process types Table 1: Optional set of identifier ranges for various high-level process categories. If following this convention, charged-current (CC) processes should have identifiers in the X00-X49 block and neutral-current (NC) processes should have them in the X50-X99 block. Negative process IDs should be used for electromagnetic interactions in event generators that include them.
In Table 1, distinct ranges for shallow inelastic scattering and deep inelastic scattering are provided because existing neutrino generators often make such a distinction in both the modeling details and simulated event metadata. It should be noted that the boundary between these two is not a priori well defined and choices varies significantly in current neutrino scattering simulations, usually defined with reference to some boundary value in the invariant mass of the hadronic final state. The lack of consensus and common definitions hinders progress in modeling neutrino-induced low energy hadronization. New efforts in this area are sorely needed and would facilitate collaboration with the collider community, as discussed in Section 2.4 of Ref. [4].
@@
The total cross section for the incoming beam particle, with its specific energy, to interact with the target particle should be stored in a HepMC3::DoubleAttribute on the HepMC3::GenEvent, named "tot_xs". See G.R.6 for how to communicate cross section units. -
@@
The total cross-section for the selected process ID for the incoming beam particle, with its specific energy, to interact with the target particle should be stored in a HepMC3::DoubleAttribute on the HepMC3::GenEvent, named "proc_xs". See G.R.6 for how to communicate cross section units. -
@@
Some simulations build up an estimate of the flux-averaged total cross section as they run, which makes implementing G.C.2 impractical in many cases. As an alternative, the built-in attribute HepMC3::GenCrossSection, accessed via GenEvent::cross_section, should be used to store the current estimate of . A user should only ever need to use the estimate provided with the last event that they read to correctly scale an event rate to a cross-section prediction. This means that statistically correct predictions can be made without reading an event vector to the end. The
HepMC3::GenCrossSection::cross_section_errors data member can be used to decide when enough events have been read to reach some desired statistical precision on the total cross section. The best estimate from the generator run will always be provided on the final event in a vector.For event generators that do not currently provide the value of in the output, Appendix A provides an algorithm for computing a running estimate, with associated Monte Carlo statistical uncertainty, as events are generated.
@@
If the "lab_pos" attribute vector contains three entries then it is assumed to only contain the spatial position of the event. If it contains four entries, then the fourth entry is interpreted as the time of the event in seconds.
2.4.3 Suggestions
@@
For more complex beam simulations that can not adequately be described by a single energy or energy histogram (see G.C.4), it is suggested that the full parent decay history is included in the HepMC3::GenEvent. A full set of conventions for the description of beam particle production and parent particle decay chains (for the case of neutrino beams) is currently outside the scope of this specification, but generator implementations can signal that they adhere to this suggestion to notify users that some or all of the beam particle production information is included in the event.
@@ 2.5 Vertex Information
The vertices in a HepMC3 event are used to connect groups of incoming and outgoing particles. For the vertex information, there is only one requirement and one convention in the present version of the NuHepMC standard.
2.5.1 Requirements
@@
The standard HepMC3 range of reserved values for HepMC3::GenVertex::status is extended to include the concept of a primary vertex, corresponding to the primary hard-scattering process (i.e., the one labeled according to E.R.3), and a final state interaction (FSI) summary vertex. The full set of defined status codes can be found in Table 2. Implementations are free to define specific vertex status codes to refer to individual FSI (or ISI) processes and output as much information as they require. However, a single summary vertex may be useful for some purposes if the full FSI history is very detailed or not often needed by users. See Figure 4 for an example of an event with such a summary vertex.Status Code Meaning Usage 0 Not defined Do not use 1 Primary interaction vertex Recommended for all cases 2 FSI Summary vertex Recommended for all cases 3-20 Reserved for future NuHepMC standards Do not use 21-999 Generator-dependent For generator usage Table 2: Set of vertex status codes Any secondary vertex included within a NuHepMC event may have a status between 21 and 999. G.R.9 requires that all generator-specific status codes must be fully described by attributes stored in the HepMC3::GenRunInfo.
2.5.2 Conventions
@@
When an interaction with a nucleon bound within a nucleus with definite kinematics is simulated, a HepMC3::GenVertex corresponding to the separation of the struck nucleon and the nuclear remnant may be included and assigned status code . If this convention is signaled via G.R.4, then status code need not be included in the implementation of G.R.9.
@@ 2.6.1 Requirements
The standard HepMC3 range of reserved values for
HepMC3::GenParticle::status is slightly extended to include the concept of a target particle. For neutrino scattering, this will usually be a target nucleus. The status codes are defined in Table 3.Status Code Description Usage 0 Not defined Do not use 1 Undecayed physical particle Recommended for all cases 2 Decayed physical particle Recommended for all cases 3 Documentation line Used for in/out particles in the primary process 4 Incoming beam particle Recommended for all cases 5-10 Reserved for future HepMC3 standards Do not use 11-19 Reserved for future NuHepMC standards Do not use 20 Target particle Recommended for all cases 21-200 Generator-dependent For generator usage 201- Simulation-dependent For simulation software usage Table 3: Particle status codes Note especially that any incoming real particle must have a status code of 4 or 20, and any outgoing real particle must have a status code of 1. Special care must be taken when including the effects of initial-state and final-state interactions.
Any internal particle included within a NuHepMC event may have a status in the range than 21-200. G.R.10 requires that all generator-specific status codes must be fully described by attributes stored in the HepMC3::GenRunInfo.
2.6.2 Conventions
@@
When an interaction with a bound nucleon with definite kinematics is simulated, the internal HepMC3::GenParticle corresponding to the bound nucleon should have status code 21. If this convention is signaled via G.R.4, then status code 21 need not be included in the implementation of G.R.10. -
@@
HepMC3 places restrictions on all external particles in the event graph to facilitate automatic checking of four momentum conservation at the event graph level. As a result, a new particle number 2009900000 is defined to identify a nuclear remnant that is not precisely handled by the interaction simulation. If a generator does correctly simulate any nuclear remnant, such that the total mass and energy are well defined, then this convention is unnecessary and the remnant can be written out as a particle with a standard PDG nuclear number and mass and momentum set as normal. A pseudoparticle with this number is explicitly an implementation detail that should be used to abide by HepMC3 constraints to not have vertices with no outgoing particles. Any pseudoparticle with this number should not be considered for physics analyses or onward simulation. The number is chosen, according to the PDG scheme [27, Sec. 45, p. 733], to be outside the range reserved for nuclear and quark-content particles. Its status as a non-standard particle code is signaled by the 6th and 7th least significant digits being set to 9.If the the kinematics of the nuclear remnant are not known, but the proton and neutron numbers are well-defined, then the corresponding PDG nuclear number may be added as a HepMC3::IntAttribute on the HepMC3::GenParticle named "remnant_particle_number".
@@ 3 Flux-averaged total cross section
Section 3.1 derives an expression for that is generally applicable to most simulations of interest for neutrino experiments, including those with a time-dependent neutrino source and a full treatment of the detector geometry. Example methods for obtaining Monte Carlo estimators of suitable for storage via E.C.4 are described in Appendix A. Section 3.2 describes some simple cases in which may be calculated directly, making them good candidates for situations in which event generators may implement G.C.2.
@@ 3.1 Derivation
is the flux-averaged total cross section. For a discussion on how these might be calculated using Monte-Carlo methods see Appendix A.
@@ 3.2 Analytic calculation under simplifying assumptions
In certain simple cases, the flux-averaged total cross section is known at the start of the run and can be computed analytically to implement G.C.2. A common example is the case of a point target of type that is exposed to uni-directional beam of (anti)neutrinos of species with a fixed emission time. In this case, we may write the differential flux and target density as
@@ 4 NuHepMC Generator Predictions
Using preliminary tools for converting proprietary neutrino event formats to the NuHepMC standard defined above, we demonstrate the utility of the common format with some cross-section predictions. See Tab. 4 for the current status of implementations of NuHepMC in the most widely-used neutrino event generators. 555An up to date table of the current status of adoption within the generators can be found in the NuHepMC Spec repository: https://github.com/NuHepMC/Spec. It is important to note that once built-in support for NuHepMC is adopted within a generator, the public converter from the previous generator-specific formats to NuHepMC will be deprecated.
| Generator | Output Formats | Public Converter |
| Achilles | NuHepMC | N/A |
| GENIE | GHep (ROOT TTree), NuHepMC in review | No |
| GiBUU | LHA XML, NuHepMC in 2025 release | N/A |
| MARLEY | NuHepMC (in upcoming v2 release) | N/A |
| NEUT | neutvect (ROOT TTree) | Yes |
| NuWro | event1 (ROOT TTree) | Yes |
A comparison to a recent neutrino-argon cross-section measurement from the MicroBooNE collaboration [29] is shown on the left-hand side of Figure 8. This comparison was made with the NUISANCE [22] framework,666The NUISANCE framework can be obtained at: https://github.com/NUISANCEMC/nuisance. which before this implementation of NuHepMC would have to have been built against GENIE, NEUT, and NuWro binaries of compatible versions to be able to generate the predictions shown in the figure. In the present workflow, results can be obtained using only generator-agnostic tools for interpreting events stored in the NuHepMC format. This reduction of technical requirements will dramatically lower the barrier to making high-quality prediction–data comparisons for the neutrino community.
It is particularly worth noting that the different generators included in this comparison provide critical cross-section scaling information in different ways. Thanks to the provisions in this standard, NUISANCE is able to use G.R.4 to determine which representation of is available and then automatically scale the selected event rate to a differential cross-section prediction without any user input or generator-specific code or knowledge. For reference, the NEUT and NuWro event distributions are scaled via G.C.2, while GiBUU, GENIE, and Achilles are scaled via E.C.4.
We want to emphasize that the comparison to data is solely a demonstration of the prediction pipeline and not meant as a physics statement on the best or most recent predictions from each of the generators for the given measurement.
@@
where is the flux-averaged total cross section (see Sec. 3), is the statistical weight of the -th event, is the number of generated events, and is the width of the -th bin. The symbol evaluates to unity when the value of from the -th event falls within the -th bin and zero otherwise. In the case of unit-weight events, the value of for all events, and . One thus recovers the traditional definition of these estimators for unweighted events.
@@
where we have assumed that is exactly known. For cases where it is estimated using Monte Carlo techniques rather than directly calculated, the statistical uncertainty discussed in Appendix A also applies.